# On the Brun–Titchmarsh theorem. I

### v3

@@ Abstract.
The classical Brun–Titchmarsh theorem gives an upper bound, which is of correct order of magnitude in the full range, for the number of primes $p\leqslant x$ satisfying $p\equiv a\,(\operatorname{mod}{q})$. We strength this inequality for $\log q/\log x$ in different ranges, improving upon previous works by Motohashi, Goldfeld, Iwaniec, Friedlander and Iwaniec, and Maynard for general or special moduli. In particular, we are able to beat Iwaniec’s barrier $q<x^{9/20-}$, and improve all the existing inequalities in the range $x^{9/20}\ll q<x^{1/2-}$ by utilizing bilinear or trilinear structures in the remainder terms of linear sieve. The proof is based on various estimates for character and exponential sums, by appealing to arithmetic exponent pairs and bilinear forms with algebraic trace functions from $\ell$-adic cohomology, trilinear forms with Kloosterman fractions, and sums of Kloosterman sums from spectral theory of automorphic forms, as well as large value theorem for Dirichlet polynomials.
@@ 1.1. Main results
for all $1\leqslant q<x,$ and this bears the so-called Brun–Titchmarsh theorem. In what follows, we assume $q\sim x^{\varpi}$ with $\varpi\in[0,1[$, and explore a positive constant $C(\varpi)$ as small as possible, such that
@@
[…]
@@
Motohashi [Mo73] took $C(\varpi)={16}/(8-3\varpi)$ for $\varpi<1/3,$ and the admissible range was shortly extended by Goldfeld [Go75] to $\varpi<24/71.$ Iwaniec [Iw82] obtained the same constant for all $\varpi<9/20,$ for which he utilized Burgess’s estimate for short character sums:
@@
holds for any non-trivial Dirichlet character modulo $q$ and $r\in\{1,2,3\}$ (see [Bu63, Bu86]). Taking $r=2$, (1.7) implies the so-called Burgess bound $L(\frac{1}{2},\chi)\ll q^{3/16+\varepsilon}$ for Dirichlet $L$-functions. The restriction $\varpi<9/20$ is quite essential in Iwaniec’s approach. On the other hand, the choice $C(\varpi)=4/(2-\varpi)$ for $\varpi\in[9/20,1/2[$ has been obtained implicitly by Iwaniec [Iw82]. This has been illustrated by Baker [Ba96]. One may compare our bounds in Theorem 1.1 with Iwaniec–Baker’s constant. In particular, we may take $C(\frac{9}{20})\approx 2.4814$ for general $q$ and $C(\frac{9}{20})\approx 2.441$ for prime $q$, however Iwaniec–Baker needed to take $C(\frac{9}{20})\approx 2.5806.$
As one may check, the choices (1.5) and (1.6) for $C^{*}(\varpi)$ are always positive for all $\varpi\in[9/20,1/2[$, and Theorem 1.1 improves all previous known bounds even if one takes $C^{*}(\varpi)=0$ in (1.4). The gain of positive choices of $C^{*}(\varpi)$ comes from a suitable Buchstab iteration in the application of sieve methods, so that one may create trilinear structures in the remainder terms with the aid of the well-factorization of the linear sieve weight constructed by Iwaniec [Iw80]. For general moduli, we estimate a twisted moment of character sums (see Lemma 2.6) by appealing to Kloostermania developed by Deshouillers and Iwaniec [DI82]; and in the prime moduli case, the above choice for $C(\varpi)$ is independent from the Ramanujan–Petersson conjecture, and it will follow from another different approach, by appealing to a recent work of Blomer, Humphries, Khan and Milinovich [BHKM20] on twisted fourth moment of Dirichlet $L$-functions to prime moduli (see Lemma 2.7 below).
@@
The statementss of Theorems 1.1 and 1.2 are a bit complicated at first glance. The following table presents our choices of $C(\varpi)$, for some specialized $\varpi>1/2,$ and how much we can improve upon Iwaniec [Iw82].
$\varpi$ Our $C(\varpi)$ Iwaniec’s $C(\varpi)$ Improvement
$16/31$ $3.3067$ $3.3514$ $1.4\%$
$12/23$ $3.3455$ $3.4074$ $1.8\%$
$32/61$ $3.3889$ $3.4366$ $1.3\%$
$8/15$ $3.4615$ $3.5294$ $1.9\%$
$7/13$ $3.5254$ $3.5862$ $1.6\%$
$6/11$ $3.5918$ $3.6667$ $2.0\%$
Table 1. Improved $C(\varpi)$
To see the strength of Theorems 1.1 and 1.2, we would like to recall some previous admissible values of $C(\varpi)$ obtained by a series of big names.
@@
]0,1[]0,1[] 0 , 1 [ existence Titchmarsh [Ti30]
]0,1[]0,1[] 0 , 1 [ $2/(1-\varpi)$ van Lint & Richert [LR65]
@@
]2/5,1/$2]$ / 5 , 1 / 2 ] $2/(2-3\varpi)$ Motohashi [Mo73]
]0,24/71[]0,24/71[] 0 , 24 / 71 [ $16/(8-3\varpi)$ Goldfeld [Go75]
]0,9/20[]0,9/20[] 0 , 9 / 20 [ $16/(8-3\varpi)$ Iwaniec [Iw82, Theorem 3]
@@
[9/20,1/2[[9/20,1/2[[ 9 / 20 , 1 / 2 [ $4/(2-\varpi)$ Baker [Ba96, Theorem 1]
[6/11,1[[6/11,1[[ 6 / 11 , 1 [ $(2-(\frac{1-\varpi}{4})^{6})/(1-\varpi)$ Friedlander & Iwaniec [FI97, Theorem 1]
$[1-\delta,1[$ $(2-c_{0}(1-\varpi)^{2})/(1-\varpi)$ Bourgain & Garaev [BG14, Theorem 4]
]0,1/8[]0,1/8[] 0 , 1 / 8 [ $2$ Maynard [Ma13]
@@
The second part of this paper is to reduce the admissible value of $C(\varpi)$ further by assuming that $q$ has good factorizations. For a small constant $\eta>0$, we say $q$ is $\eta$-smooth if all prime factors of $q$ do not exceed $q^{\eta}.$ The smoothness assumption allows us to utilize the factorization of $q$ to control the resultant algebraic exponential sums very effectively. In particular, one may go far beyond the so-called Pólya–Vinogradov barrier for short exponential and character sums, as a consequence of the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs developed by Wu, Xi and Sawin [WX21]. One is referred to Section 3.2 for details.
The following theorem shows that we may go further when $\varpi$ is close to $2/3$, in particular, larger than $2/3$ in a explicit way, provided that $q$ is squarefree and smooth enough.
@@
[…]
Theorem 1.3 is in fact a special case of the following slightly general formulation on taking $(\kappa,\lambda)=(1/6,2/3)$.
@@
Theorems 1.3 and 1.4 are stronger than van Lint and Richert [LR65] and Iwaniec [Iw82] when $\varpi$ is around $2/3$ Put
@@
To have $f(\kappa,\lambda)$ or $g_{\varpi}(\kappa,\lambda)$ as large as possible, it requires to choose $(\kappa,\lambda)$ by applying $A$- and $B$-processes to $(0,1)$ suitably; we refer to Section 3.2 for the terminology and conventions. The optimization process can be conducted following the algorithm in [GK91, Section 5.3]. In particular, for $\varpi=2/3,$ we find
@@
and it suffices to minimize $\kappa+\lambda.$ This task matches exactly in determining Rankin’s constant for Riemann zeta functions. According to Rankin (see also [GK91, Section 5.4]), the minimal value for $\kappa+\lambda$ should be $0.829021\cdots$ by choosing
@@
in which case we get $g_{\varpi}(\kappa,\lambda)=0.723659\cdots$. This yields $C(2/3)=4/g_{\varpi}(\kappa,\lambda)\approx 5.52746$ in Theorem 1.4. Note that van Lint and Richert [LR65] obtained the constant $62/(1-\varpi)=$ 6 for $\varpi=2/3.$
@@
in ($,$ provided that $\eta$ is small enough.
Theorem 1.5 can be compared with the following conditional result due to Motohashi [Mo74]. Under the Linderlöf Hypothesis for Dirichlet $L$-functions, he proved that
$C(\varpi)=\begin{cases}2,\quad&\varpi\in[0,~{}1/3],\\ 2/(2-3\varpi),&\varpi\in[1/3,~{}1/2],\end{cases}$
is admissible in (1.3).
@@
1.2. Ingredients of proof and comparisons with previous works
Before discussing the ingredients of our proof, it would be better to give a short survey on the methods employed in previous works.
  • Both of van Lint and Richert [LR65] and Motohashi [Mo74, Mo73] utilized Selberg’s sieve, which is more convenient than Brun’s sieve at that time. Recall that one takes $C(\varpi)=2/(1-\varpi)$ in [LR65], and the constant $2$ is resulted naturally by applications of the linear sieve with its best possible main term. The above denominator $(1-\varpi)$ comes from the level $(x/q)^{1-\varepsilon}$ after a trivial treatment of error terms. It was Motohashi [Mo74, Mo73] who first dealt with the error term in a non-trivial manner, and he successfully employed the bilinear structure in Selberg’s sieve. He employed multiplicative characters, and a basic tool is the mean value of Dirichlet $L$-functions, which prevents him from improving van Lint and Richert [LR65] for $\varpi>1/2.$
@@ 1.2. Ingredients of proof and comparisons with previous works
    holds with $c=2.8$, for all $q<x/c_{0}$ with some $c_{0}>1.$ Selberg employed his own $\Lambda^{2}$ sieve, and published this inequality in [Se91, Page 233]. In another scene, Montgomery and Vaughan [MV73] developed the large sieve, from which they deduced that (1.8) is valid for all $1\leqslant q<x$ and $c=0$.
  • Iwaniec [Iw82] developed a new linear sieve of combinatorial type with well-factorable weights. This far-reaching work was originally motivated by Chen [Ch75] and Motohashi [Mo74], and the merit is that one may optimize the balances between convolutions in the error term, and the sign changes among them can be captured more effectively. In particular, this well-factorization in the error term allows Iwaniec [Iw82] to employ (Weil’s bound for) Kloosterman sums to beat the barrier $\varpi\approx 1/2$. He was able to beat van Lint and Richert [LR65] even if $\varpi$ is close to $2/3-.$ On the other hand, it is worthwhile to mention that Motohashi [Mo99] showed another bilinear structure in the Selberg $\Lambda^{2}$ sieve, which has a similar flavor to that in [Iw82]. In particular, he proved that $C(\varpi)=16/(8-3\varpi)$ is admissible for all $\varpi<9/20,$ which recovered the work of Iwaniec.
@@
    where $m$ and $n$, as prime variables, are restricted to very special but short segments. In particular, they succeed in improving van Lint and Richert [LR65] for all $\varpi\in[2/3,1[.$ To apply Karatsuba’s estimate for (1.9), Friedlander and Iwaniec [FI97] introduced Buchstab’s iteration so that one may create prime variables in the sifting functions. By employing ideas from additive combinatorics (sum-product phenomenon), Bourgain and Garaev [BG14] can bound (1.9) without restricting $m,n$ to primes. Therefore, Buchstab’s iteration is not necessary, and they can give a further improvement over [FI97].
@@
with $MN$ as large as possible, where $MNL=x.$ The coefficient $-β\boldsymbol{\alpha}*\bbeta$ comes from the linear sieve of Iwaniec [Iw82], and each of $-β\boldsymbol{\alpha},\bbeta$ is bounded by divisor functions. There are at least two approaches to detect the congruence condition $mnl\equiv a\,(\operatorname{mod}{q}),$ via multiplicative (Dirichlet) characters or additive characters.
@@
One may easily see this approach does not work unless $\varpi<1/2,$ by expecting that there exist (at most) squareroot cancellations on the left hand side. After suitable applications of Hölder’s inequality, one may arrive at bounding the twisted moment
$\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{r\sim R}c_{r}\chi(r)\Big{|}^{2}\Big{|}\sum_{l\sim L}\chi(l)\Big{|}^{2k}.$
The underlying idea is to prove squareroot cancellations among the averages over $r$ and $l$. This expectation cannot be realized when $R$ or $k$ is suitably large, hence it becomes necessary to control large values of Dirichlet polynomials.
In his original proof, Iwaniec [Iw82] applied fourth moment of short character sums or Dirichlet $L$-functions (i.e., $k=2$ and $R=1$). As an approximation to the sixth moment, we may also allow, in the case $k=2$, the length $R$ to be suitably large so that a kind of Linderlöf Hypothesis on average can be achieved. This will appeal to the recent work of Blomer, Humphries, Khan and Milinovich [BHKM20], in which $q$ is restricted to primes and $R$ can be as large as $q^{1/4}$.
In another direction, the coefficient $(c_{r})$ has a convolution structure, coming from Iwaniec’s new construction of the Rosser–Iwaniec sieve. We utilize this special structure of $(c_{r})$ as efficiently as possible, and for $k=1$ we may give another kind of Linderlöf Hypothesis on average for $R\approx q^{(5-2\vartheta)/8}$. Together with large sieves (Lemma 2.1), this can also produce a new level of distribution in linear sieves. The advantage is that $q$ is not necessarily a prime, but we are restricted by the possibility of exceptional spectrum of the non-Eucliean Laplacian. It is also worthwhile to remark that if we may factorize the coefficient $(c_{r})$ as we wish, this argument can produce the same level without appealing to the work of Blomer, Humphries, Khan and Milinovich [BHKM20], modulo the Ramanujan–Petersson conjecture.
@@
(2) If $q$ has no good factorizations, say $q$ is a large prime, we also employ the Pólya–Vinogradov method, but do not go with Weil’s bound directly. Alternatively, we take the advantages of averages over $h_{1},h_{2},n_{1},n_{2}$, as well as the dual sum resulted from the Pólya–Vinogradov method. By grouping $h_{1},h_{2},n_{1},n_{2}$ in certain ways, we may create some bilinear and trilinear forms with Kloosterman sums. This produces possibilities to capture sign changes of Kloosterman sums. For the bilinear form, we need to control a sum of product of many Kloosterman sums, which will appeal to [FGKM14] or [Xi17]; while for the trilinear form, we reduce the problem to bounding the bilinear form

### v4

@@ Abstract.
The classical Brun–Titchmarsh theorem gives an upper bound, which is of correct order of magnitude in the full range, for the number of primes $p\leqslant x$ satisfying $p\equiv a\,(\operatorname{mod}{q})$. We strengthen this inequality for different ranges of $\log q/\log x$, improving upon previous works by Motohashi, Goldfeld, Iwaniec, Friedlander and Iwaniec, and Maynard for general or special moduli. In particular, we are able to beat Iwaniec’s barrier $q<x^{9/20-}$, and improve all existing inequalities in the range $x^{9/20}\ll q<x^{1/2-}$ by utilizing bilinear or trilinear structures in the remainder terms of linear sieve. The proof is based on various estimates for character and exponential sums, which we derive by appealing to arithmetic exponent pairs and bilinear forms with algebraic trace functions from $\ell$-adic cohomology, trilinear forms with Kloosterman fractions, and sums of Kloosterman sums from spectral theory of automorphic forms, as well as large value theorem for Dirichlet polynomials.
@@ 1.1. Main results
for all $1\leqslant q<x,$ and this bears the name Brun–Titchmarsh theorem. In what follows, we assume $q\sim x^{\varpi}$ with $\varpi\in[0,1[$, and seek a positive constant $C(\varpi)$, as small as possible, such that
@@
in (1.3) with some constant $C^{*}(\varpi)\geqslant 0.$ In general, we may choose
@@
It is worthwhile to make a remark on the merit of Theorem 1.1. Note that the choices (1.5) and (1.6) for $C^{*}(\varpi)$ are always positive for all $\varpi\in[9/20,1/2[$, and Theorem 1.1 improves all previously known bounds even if taking $C^{*}(\varpi)=0$ in (1.4). The gain coming from positive choices of $C^{*}(\varpi)$ arises from a suitable Buchstab iteration in the application of sieves, so that one may create trilinear structures in the remainder terms with the aid of the well-factorization of the linear sieve weight constructed by Iwaniec [Iw80]. For general moduli, we estimate a twisted moment of character sums (see Lemma 2.6) by appealing to Kloostermania developed by Deshouillers and Iwaniec [DI82]; and in the prime moduli case, the above choice for $C(\varpi)$ is independent of the Ramanujan–Petersson conjecture, and it will follow from another different approach, by appealing to a recent work of Blomer, Humphries, Khan and Milinovich [BHKM20] on twisted fourth moment of Dirichlet $L$-functions to prime moduli (see Lemma 2.7 below).
In fact, Motohashi [Mo73] took $C(\varpi)={16}/(8-3\varpi)$ for $\varpi<1/3,$ and the admissible range was later extended by Goldfeld [Go75] to $\varpi<24/71.$ Iwaniec [Iw82] obtained the same constant for all $\varpi<9/20,$ for which he utilized Burgess’s estimate for short character sums:
@@
holds for any non-trivial Dirichlet character modulo $q$ and $r\in\{1,2,3\}$ (see [Bu63, Bu86]). Taking $r=2$, (1.7) implies the so-called Burgess bound $L(\frac{1}{2},\chi)\ll q^{3/16+\varepsilon}$ for Dirichlet $L$-functions. The restriction $\varpi<9/20$ is quite essential in Iwaniec’s approach. On the other hand, the choice $C(\varpi)=4/(2-\varpi)$ for $\varpi\in[9/20,1/2[$ has been obtained implicitly by Iwaniec [Iw82]. This has been illustrated by Baker [Ba96]. One may compare our bounds in Theorem 1.1 with Iwaniec–Baker’s constant. In particular, we may take $C(\frac{9}{20})\approx 2.4814$ for general $q$ and $C(\frac{9}{20})\approx 2.441$ for prime $q$, whereas Iwaniec–Baker had to take $C(\frac{9}{20})\approx 2.5806.$
@@
To see the strength of Theorems 1.1 and 1.2, we recall some previous admissible values of $C(\varpi)$ obtained by a series of big names.
@@
]0,1[]0,1[ existence Titchmarsh [Ti30]
]0,1[]0,1[ $2/(1-\varpi)$ van Lint & Richert [LR65]
@@
]2/5,1/2]]2/5,1/2] $2/(2-3\varpi)$ Motohashi [Mo73]
]0,24/71[]0,24/71[ $16/(8-3\varpi)$ Goldfeld [Go75]
]0,9/20[]0,9/20[ $16/(8-3\varpi)$ Iwaniec [Iw82, Theorem 3]
@@
$[9/20,1/2[$ $4/(2-\varpi)$ Baker [Ba96, Theorem 1]
$[6/11,1[$ $(2-(\frac{1-\varpi}{4})^{6})/(1-\varpi)$ Friedlander & Iwaniec [FI97, Theorem 1]
$1-\delta,1[$ $(2-c_{0}(1-\varpi)^{2})/(1-\varpi)$ Bourgain & Garaev [BG14, Theorem 4]
]0,1/8[]0,1/8[ 22 Maynard [Ma13]
@@
The second aim of this paper is to reduce the admissible value of $C(\varpi)$ further by assuming that $q$ has good factorizations. For a small constant $\eta>0$, we say $q$ is $\eta$-smooth if all prime factors of $q$ do not exceed $q^{\eta}.$ The smoothness assumption allows us to utilize the factorization of $q$ to control the resultant algebraic exponential sums very effectively. In particular, one may go far beyond the so-called Pólya–Vinogradov barrier for short exponential and character sums, as a consequence of the general form of the $q$-analogue of van der Corput method and arithmetic exponent pairs developed by Wu, Xi and Sawin [WX21]. We refer the reader to Section 3.2 for details.
The following theorem shows that we may go further when $\varpi$ is close to $2/3$, and in particular for explicit ranges with $\varpi>2/3$, provided that $q$ is squarefree and sufficiently smooth.
@@
in (1.3) for all $\varpi\in[3/10,3/4],$ provided that $\eta$ is small enough.
Theorem 1.3 is in fact a special case of the following general formulation with $(\kappa,\lambda)=(1/6,2/3)$.
@@
Theorems 1.3 and 1.4 are stronger than the results of van Lint and Richert [LR65] and that of Iwaniec [Iw82] when $\varpi$ is around $2/3$, assuming that $q$ is $\eta$-smooth and squarefree. Put
@@
To have $f(\kappa,\lambda)$ or $g_{\varpi}(\kappa,\lambda)$ as large as possible, one needs to choose $(\kappa,\lambda)$ by applying $A$- and $B$-processes to $(0,1)$ suitably; we refer to Section 3.2 for the terminology and conventions. The optimization process can be conducted following the algorithm in [GK91, Section 5.3]. In particular, for $\varpi=2/3,$ we find
@@
and it suffices to minimize $\kappa+\lambda.$ This task matches exactly the problem of determining Rankin’s constant for Riemann zeta function. According to Rankin (see also [GK91, Section 5.4]), the minimal value for $\kappa+\lambda$ should be $0.829021\cdots$ by choosing
@@
in which case we get $g_{\varpi}(\kappa,\lambda)=0.723659\cdots$. This yields $C(2/3)=4/g_{\varpi}(\kappa,\lambda)\approx 5.52746$ in Theorem 1.4. Note that van Lint and Richert [LR65] obtained the constant $2/(1-\varpi)=6$ for $\varpi=2/3.$
@@
[…]
Theorem 1.5 can be compared with the following conditional result due to Motohashi [Mo74]. He proved that the Lindelöf Hypothesis for Dirichlet $L$-functions allows one to take in (1.3)
$C(\varpi)=\begin{cases}2,\quad&\varpi\in[0,~1/3],\\ 2/(2-3\varpi),&\varpi\in[1/3,~1/2].\end{cases}$
1.2. Ingredients of proof and comparisons with previous works
The statements of Theorems 1.1 and 1.2 are a bit complicated at first glance. The following table presents our choices of $C(\varpi)$, for some specific $\varpi>1/2,$ and shows how much we can improve upon Iwaniec [Iw82]. It is worthwhile to record that any small improvements on the constants rely heavily on various estimates for exponential and character sums, individually or on average. More precisely, we appeal to spectral theory of automorphic forms for sums of Kloosterman sums, and to $\ell$-adic cohomology for short character sums and algebraic exponential sums.
$\varpi$ Our $C(\varpi)$ Iwaniec’s $C(\varpi)$ Improvement
$16/31$ $3.3067$ $3.3514$ $1.4\%$
$12/23$ $3.3455$ $3.4074$ $1.8\%$
$32/61$ $3.3889$ $3.4366$ $1.3\%$
$8/15$ $3.4615$ $3.5294$ $1.9\%$
$7/13$ $3.5254$ $3.5862$ $1.6\%$
$6/11$ $3.5918$ $3.6667$ $2.0\%$
Table 2. Improved $C(\varpi)$
@@ 1.2. Ingredients of proof and comparisons with previous works
Before discussing the ingredients of our proof, it would be better to give a short survey of the methods employed in previous works.
  • Both van Lint and Richert [LR65] and Motohashi [Mo74, Mo73] utilized Selberg’s sieve, which is more convenient than Brun’s sieve at that time. Recall that one takes $C(\varpi)=2/(1-\varpi)$ in [LR65], and the constant $2$ resulted from applications of the linear sieve with its best possible main term. The above denominator $(1-\varpi)$ comes from the level $(x/q)^{1-\varepsilon}$ after a trivial treatment of error terms. It was Motohashi [Mo74, Mo73] who first dealt with the error term in a non-trivial manner, and he successfully employed the bilinear structure in Selberg’s sieve. He employed multiplicative characters, and a basic tool in his work was the mean value of Dirichlet $L$-functions. Such an approach did not allow him to improve upon the result of van Lint and Richert [LR65] for $\varpi>1/2$.
@@
    holds with $c=2.8$, for all $q<x/c_{0}$ with some $c_{0}>1.$ Selberg employed his own $\Lambda^{2}$ sieve, and published this inequality in [Se91, Page 233]. In an another direction, Montgomery and Vaughan [MV73] developed the large sieve, from which they deduced that (1.8) is valid for all $1\leqslant q<x$ and $c=0$.
  • Iwaniec [Iw82] developed a new linear sieve of combinatorial type with well-factorable weights. This far-reaching work was originally motivated by Chen [Ch75] and Motohashi [Mo74], and the merit is that one may optimize the balances between convolutions in the error term, and the sign changes among them can be captured more effectively. In particular, this well-factorization in the error term allows Iwaniec [Iw82] to employ (Weil’s bound for) Kloosterman sums to beat the barrier $\varpi\approx 1/2$. He was able to beat van Lint and Richert [LR65] even if $\varpi$ is close to 2/3.2/3-. On the other hand, it is worthwhile to mention that Motohashi [Mo99] showed another bilinear structure in the Selberg $\Lambda^{2}$ sieve, which has a similar flavor to that in [Iw82]. In particular, he proved that $C(\varpi)=16/(8-3\varpi)$ is admissible for all $\varpi<9/20,$ which recovered the work of Iwaniec.
@@
    where $m$ and $n$, as prime variables, are restricted to very special but short segments. In particular, they succeeded in improving van Lint and Richert [LR65] for all $\varpi\in[2/3,1[.$ To apply Karatsuba’s estimate for (1.9), Friedlander and Iwaniec [FI97] introduced Buchstab’s iteration so that one may create prime variables in the sifting functions. By employing ideas from additive combinatorics (sum-product phenomenon), Bourgain and Garaev [BG14] can bound (1.9) without restricting $m,n$ to primes. Therefore, Buchstab’s iteration is not necessary, and they can give a further improvement over [FI97].
@@
[…]
@@
One may easily see this approach does not work unless $\varpi<1/2,$ by expecting that there exist (at most) square-root cancellations on the left hand side. After suitable applications of Hölder’s inequality, one may arrive at bounding the twisted moment
(1.10) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big|\sum_{r\sim R}c_{r}\chi(r)\Big|^{2}\Big|\sum_{l\sim L}\chi(l)\Big|^{2k}.$
The underlying idea is to prove square-root cancellations among the averages over $r$ and $l$. This expectation cannot be realized when $R$ or $k$ is suitably large, hence it becomes necessary to control large values of Dirichlet polynomials.
In his original proof, Iwaniec [Iw82] studied the fourth moment of character sums twisted by Dirichlet polynomials (i.e., $k=2$ in (1.10)). In contrast with Iwaniec’s original argument, one of the novelties in proving Theorem 1.1 is that we use the Buchstab identity in the implementation of linear sieves. As a consequence, we are led to study two types of second moments of character sums twisted by Dirichlet polynomials (i.e., $k=1$ in (1.10)). For the first type, the coefficient $(c_{r})$ is arbitrary, and we appeal to the work of Bettin and Chandee [BC18] on trilinear forms involving Kloosterman fractions. For the second type, the coefficient $(c_{r})$ has a convolution structure, arising from the Buchstab identity and Iwaniec’s new construction of the Rosser–Iwaniec sieve. We exploit this special structure of $(c_{r})$ as efficiently as possible: it allows us to apply estimates for sums of Kloosterman sums coming from spectral theory of automorphic forms, and in this way we derive a kind of Lindelöf Hypothesis on average for $R\approx q^{(5-2\vartheta)/8}$ (for further discussions, see Section 8).
By employing Hölder’s inequality in a different way, we are instead led to bound fourth moment of Dirichlet $L$-functions twisted by Dirichlet polynomials (intimately related to $k=2$ in (1.10)) rather than twisted second moments of character sums described above. In this setting we can apply the recent work of Blomer, Humphries, Khan and Milinovich [BHKM20], where $q$ is restricted to primes and $R$ can be as large as $q^{1/4}$, and thereby obtain another kind of Lindelöf Hypothesis on average. This serves as a key ingredient in establishing the second assertion of Theorem 1.1.
We remark that the advantage of applying estimates for sums of Kloosterman sums (to twisted second moment of character sums) is that $q$ is not necessarily a prime, but we are restricted by the possible presence of exceptional spectrum of the non-Euclidean Laplacian. It is also worthwhile to remark that if we may factorize the coefficient $(c_{r})$ as we wish, this argument can produce the same level without appealing to the work of Blomer, Humphries, Khan and Milinovich [BHKM20], modulo the Ramanujan–Petersson conjecture.
An alternative approach to study (1.10) is to estimate the character sum individually. In particular, when $q$ is squarefree and has only very small prime factors, the method of arithmetic exponent pairs in [WX21] can guarantee that the sum over $l$ in (1.10) is $o(L)$ for very small $L$ and each $\chi\neq\chi_{0}$. This, amongst many other things, allows us to show that $C(\varpi)=2$ is admissible in (1.3) for $\varpi\in[1/8,~5/12[$. One will see this constant is best possible if assuming the existence of Landau–Siegel zeros.
@@
(2) If $q$ has no good factorizations, say $q$ is a large prime, we also employ the Pólya–Vinogradov method, but do not go with Weil’s bound directly. Alternatively, we take advantage of averages over $h_{1},h_{2},n_{1},n_{2}$, as well as the dual sum resulted from the Pólya–Vinogradov method. By grouping $h_{1},h_{2},n_{1},n_{2}$ in certain ways, we may create some bilinear and trilinear forms with Kloosterman sums. This produces possibilities to capture sign changes of Kloosterman sums. For the bilinear form, we need to control a sum of product of many Kloosterman sums, which will appeal to [FGKM14] or [Xi17]; while for the trilinear form, we reduce the problem to bounding the bilinear form